OpenStudy (clarence):

The bass of solid S is the region enclosed by the parabola y = 64 - 25x^2 and the x-axis. Cross-sections perpendicular to the y-axis are squares. Find the volume of the described solid S. Where do I even start?!

OpenStudy (irishboy123):

I'm on mobile so can't draw but all they're really saying is that the height of the solid z = f(x,y) =64-25x^2 So both width and height of a cross section are driven by the value of x Integrate f(x,y) over the region.

OpenStudy (irishboy123):

Actually no need for double integral. Just create a function A(x) for area at any x and integrate that dx.

OpenStudy (clarence):

So integrate 64-25x^2? @irishboy123

OpenStudy (irishboy123):

\(Width = 64-25x^2\) \(Height = 64-25x^2\) \(A(x) = W \cdot H = (64-25x^2)^2\)

OpenStudy (clarence):

Oh really? I got area to be \[\frac{ 4(64-x) }{ 25 }\]

OpenStudy (clarence):

And then integrating that from 0 to 1 to get 254/25 as my answer?

OpenStudy (clarence):

Sorry, x was meant to be y, my bad.

OpenStudy (irishboy123):



OpenStudy (clarence):

I followed my friend's suggestion and had a look at this link: http://slader.com/textbook/9780538498845-stewart-calculus-international-edition-7th-edition/362/exercises/58/

OpenStudy (irishboy123):

drat "perpendicular to" y axis means this.....  \(x = \sqrt{\dfrac{64-y}{25}}\) \(A(y) = \left[ \sqrt{\dfrac{64-y}{25}} \right]^2 \) \(V = \int\limits_{0}^{64} \; \dfrac{64-y}{25} \; dy = \dfrac{2048}{25}\)